# envname-P envname#1

### v1

@@ 3.1. Problem definition
The inputs to the problem are a set of propositional formulas $\Gamma=\{A_{1},\ldots,A_{n}\}$ representing assumptions, a propositional variable $c$ representing a conclusion, and a function $C\colon\Gamma\to\mathbb{N}$ that assigns a cost to making each assumption. The desired output is the set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of assumptions in $\Delta$ is minimized, subject to the constraint $\Delta\models c$.
Here is a simple example of the minimum trust which continues the example from Section˜2. Suppose that aho_corasick depends on memchr, and that aho_corasick was developed by Alice, while memchr was developed by Bob and has no dependencies.
@@
Table 1. Candidate Assumptions Table
A set of candidate assumptions a user can make about the safety of the aho_corasick library and their associated costs are shown in Table˜1. The solution to the minimum trust problem is the set of assumptions that entail the safety of the aho_corasick library ($c$) with minimum cost. In this example, the minimum cost necessary to show $c$ is 45, incurred from assumptions $a$ (cost 20), $b$ (cost 5), $(a\land m)\implies c$ (cost 10), and $b\implies m$ (cost 10).
@@ 3.2. Naive algorithm
Algorithm˜1 shows a naive solution to solve the minimum trust problem using an SMT solver by encoding the constraints as sentences over linear integer arithmetic and using quantifiers over propositional variables. To associate each assumption being made to a cost, we use auxiliary propositional variables $a_{1},\ldots,a_{n}$ to track each assumption $A_{1},\ldots,A_{n}$. Each $a_{i}$ is assigned to true if and only if $A_{i}\in\Delta$; in the running example, we would have $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, etc. We then encode the cost of the assumptions in $\Delta$ as a sum of if-then-else (iteite) terms; in the running example, the cost of any particular solution is given by $\texttt{cost}(a_{1},\ldots,a_{8})=ite(a_{1},100,0)+\dots+ite(a_{8},10,0)$. The constraint that the conclusion $c$ follows from the assumptions being made, $\Delta\models c$, is encoded as a formula $\texttt{valid}(a_{1},\ldots,a_{n})$; in the running example, $\texttt{valid}(a_{1},\ldots,a_{8})=\neg\exists c,m,d,a,b\>((a_{1}\implies c)\wedge\dots\wedge(a_{8}\implies(b\implies m))\wedge\neg c)$. Finally, we encode using quantifiers the fact that $\texttt{cost}(a_{1},\ldots,a_{n})$ is minimized subject to the constraint $\texttt{valid}(a_{1},\ldots,a_{n})$. In the running example, the SMT solver would produce a model that assigns $a_{5},a_{6},a_{7},a_{8}$ to true and $a_{1},a_{2},a_{3},a_{4}$ to false, incurring a cost of $45$.
@@
7:   $\Gamma\leftarrow\texttt{get\_all\_assumptions(lib)}$
@@
9:     $F\leftarrow F\wedge(a_{i_{\texttt{lib}}}\implies A_{i})$
@@
11:     $F\leftarrow F\wedge\texttt{JoinAssumptions}(\texttt{dep})$ $\triangleright$ JoinAssumptions is called recursively on each dependency
@@
13: procedure SolveMinimumTrust
14:   $F(a_{1},\ldots,a_{n})\leftarrow\texttt{JoinAssumptions(library)}$
15:   $\{v_{1},\ldots,v_{m}\}\leftarrow\texttt{atoms}(F)\setminus\{a_{1},\ldots,a_{n}\}$ $\triangleright$ atoms returns the set of propositional variables in a formula
16:   $\texttt{valid}(a_{1},\ldots,a_{n})\leftarrow\neg\exists c\exists v_{1}\ldots\exists v_{m}\>(F(a_{1},\ldots,a_{n})\wedge\neg c)$ $\triangleright$ $c$ is the conclusion ‘the library is safe’
17:   $\texttt{cost}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
18:   $\texttt{minimal}\leftarrow\forall a_{1}\ldots\forall a_{n}\>(\texttt{valid}(a_{1},\ldots,a_{n})\implies\texttt{cost}(a^{*}_{1},\ldots,a^{*}_{n})\leq\texttt{cost}(a_{1},\ldots,a_{n}))$
19:   $\texttt{S.solve}(\texttt{valid}(a^{*}_{1},\ldots,a^{*}_{n})\land\texttt{minimal})$ $\triangleright$ $(a^{*}_{1},\ldots,a^{*}_{n})$ is the solution to the minimum trust problem
@@
Algorithm˜1 recursively calls JoinAssumptions on each library in the dependency tree on line 11 to construct the formula $F(a_{1},\ldots,a_{n})$ that represents the assumptions being made, then uses the above encoding to create the constraint $\texttt{valid}(a_{1},\ldots,a_{n})$. The algorithm encodes that the cost of assumptions is minimized subject to the constraint $\texttt{valid}(a_{1},\ldots,a_{n})$ on line 18 using quantifiers. ∎
@@ 3.3. Horn clause algorithm
To address the complexity limitations of the first algorithm, we propose a second. The key idea of Algorithm˜2 is to restrict the set of assumptions to be acyclic definite Horn clauses; by doing this, the problem can be solved efficiently with a binary search reduction using quantifier-free queries to a SMT solver. The set of assumptions that can be expressed as acyclic definite Horn clauses is sufficient for all of the assumptions we identify based on prior work (Table˜2(a)).
In the running example, we can observe that the set of conjuncts of $F$ that establish the safety of the aho_corasick library ($c$) are $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, and $a_{7}\implies((a\land m)\implies c)$. Rewritten as Horn clauses with a body and a head, these conjuncts are $a_{1}\implies c$, $(a_{4}\land d)\implies c$, and $(a_{7}\land a\land m)\implies c$. Therefore, with the goal of showing that $c$ is true, the formula $a_{1}\lor(a_{4}\land d)\lor(a_{7}\land a\land m)$ must evaluate to true. The only conjunct of $F$ that establishes $d$ is $a_{3}\implies d$. Therefore, we can substitute $d$ with $a_{3}$ in the formula, resulting in $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a\land m)$. By repeating this unfolding process, we can simplify the formula to $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a_{5}\land a_{6}\land a_{8})$. This formula will evaluate to true if and only if the assumptions made entail the conclusion $c$. Algorithm˜2 uses this observation to solve the minimum trust problem for a certain library when the assumptions are acyclic definite Horn clauses. Due to binary search, Algorithm˜2 produces a logarithmic number of calls to the SMT solver with respect to the maximum cost, each with a quantifier-free formula.
@@
5:   $F\leftarrow\texttt{simplify}(F)$ $\triangleright$ simplify the formula into a conjunction of Horn clauses
6:   $\Pi\leftarrow\texttt{clauses}(F)$ $\triangleright$ get the set of Horn clauses in $F$
@@
9:     if $\texttt{head}(\pi_{i})$ is the conclusion $c$ then
10:      $E\leftarrow E\lor\texttt{body}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $c$
11:   $\leftarrow\texttt{atoms}(F)\setminus\{a_{1},\ldots,a_{n}\} \triangleright VV$ stores the set of propositional variables in $F$ other than the $a_{i}$s
@@
15:      if $\texttt{head}(\pi_{i})$ is $v_{i}$ then
16:       $T\leftarrow T\lor\texttt{body}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $v_{i}$
@@
21: procedure SolveHornMinimumTrust
22:   $F(a_{1},\ldots,a_{n})\leftarrow\texttt{JoinAssumptions(library)}$ $\triangleright$ same as in Algorithm˜1
23:   $\texttt{valid}(a_{1},\ldots,a_{n})\leftarrow\texttt{Unfold}(F)$
24:   $\texttt{cost}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
25:   $\texttt{left}\leftarrow 0$ $\triangleright$ initialize the min cost as 0
26:   $\texttt{right}\leftarrow\sum_{i=1}^{n}C(a_{i})$ $\triangleright$ the max cost is the sum of the costs of all assumptions
27:   while $\texttt{left}<\texttt{right}$ do
28:     $k\leftarrow\lfloor(\texttt{left}+\texttt{right}+1)/2\rfloor$ $\triangleright$ binary search for min cost
29:     $\texttt{cost\_leq\_k}\leftarrow\texttt{S.solve}(\texttt{valid}(a_{1},\ldots,a_{n})\wedge\texttt{cost}(a_{1},\ldots,a_{n})\leq k)$
30:     if cost_leq_k is true then
31:      $\texttt{left}\leftarrow k$
@@
33:      $\texttt{right}\leftarrow k-1$
34:   $\texttt{min\_cost}\leftarrow\texttt{left}$
@@
Algorithm˜2 is correct for assumptions consisting of acyclic definite Horn clauses.
@@
Algorithm˜2 first constructs $E$ as the disjunction of all the bodies of the Horn clauses whose heads are the conclusion $c$ (line 10). The conclusion is provable from the assumptions if and only if $E$ evaluates to true. Algorithm˜2 then proceeds to iterate through $V$ with the goal of eliminating all the propositional variables in $V$ from $E$. For each propositional variable $v_{i}$ in $V$, the algorithm constructs a formula $T$ consisting of the disjunction of the bodies of the Horn clauses whose heads are $v_{i}$ and then eliminates $v_{i}$ by substituting $T$ in place of $v_{i}$ in $E$, similar to the construction of $E$ on line 10. After the substitution, $E$ will still be a necessary and sufficient condition for the conclusion to be true. The substitution is then applied to each remaining clause in line 19; since the Horn clauses are acyclic, eliminating $v_{i}$ is guaranteed to remove $v_{i}$ from all future substitutions. This process is repeated until all propositional variables in $V$ are eliminated from $E$. Each step in Unfold maintains that $E$ is necessary and sufficient for the conclusion to be true. The formula generated by Unfold is used in a binary search to repeatedly query S for valid minimum trust solutions with different “cost budgets” until the minimum cost is found. ∎
It should be noted that the unfolding procedure in Algorithm˜2 is similar to the well-known SLD resolution algorithm for logic programming (kowalski1974predicate, ). Unfortunately, the generated formula has an exponential number of clauses in the worst case (with respect to the number of crates in the dependency tree). This formula, however, is in disjunctive normal form (after minor simplification), which means the SMT solver can solve each clause independently. Note that it is also possible to have a reduction which produces a polynomial-size formula, though this is not the reduction we use in our implementation.
@@ 3.4. Complexity
To close the section, we establish the complexity of the minimum trust problem considered as a decision problem. Define MINTRUST as follows: given a set of assumptions $\Gamma$, a conclusion $c$, and a cost function $C$, is there a set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of the assumptions in $\Delta$ is less than or equal to $k$ and $\Delta\models c$? Similarly, define HORN-MINTRUST to be the restriction of this problem where $\Gamma$ consists of Horn clauses. The following theorems establish that the unrestricted minimum trust problem is computationally infeasible (Theorem˜4) and that the Horn-restricted minimum trust problem is intractable but still within the reach of SMT solvers (Theorem˜5).
@@
We reduce from the SHORTEST-IMPLICANT-CORE (SIC) problem, which is known to be $\Sigma_{2}^{P}$-hard (schaefer2001phcompleteness, ) (umans2001implicants, ). The SIC problem asks: given a DNF formula $\varphi$, an implicant $I$ of $\varphi$ (a set of literals whose conjunction implies $\varphi$), and an integer $k$, does an implicant $I^{\prime}\subseteq I$ of size at most $k$ exist? Given an instance of SIC with a DNF formula $\varphi=\varphi_{1}\lor\dots\lor\varphi_{n}$, an implicant $I$ of $\varphi$, and an integer $k$, we construct an instance of MINTRUST with a set of assumptions $\Gamma$, a conclusion $c$, a cost function $C$, and an integer $k$. Let $\Gamma$ contain an assumption for each literal in $I$, with a cost of 1, and the assumptions $\varphi_{i}\implies c$ for each disjunct $\varphi_{i}$, with a cost of 0.
By construction, $\Delta\models c$ if and only if for some disjunct $j$, $\Delta\models\varphi_{j}$ and $(\varphi_{j}\implies c)\in\Delta$. Furthermore, $\Delta\models\varphi_{j}$ if and only if $\Delta$ contains an implicant $I^{\prime}$ of $\varphi_{j}$ such that $I^{\prime}\subseteq I$. Then, the cost associated with $\Delta$ is equal to the size of $I^{\prime}$, so $\varphi_{j}$ (and $\varphi$ by extension) has an implicant of size at most $k$ if and only if the MINTRUST instance has a solution of cost at most $k$. Therefore, an implicant $I^{\prime}\subseteq I$ of $\varphi$ of size at most $k$ exists if and only if this MINTRUST instance has a solution of cost at most $k$. ∎
@@
HORN-MINTRUST is NP-hard. We reduce from VERTEX-COVER (karp1972reducibility, ). Given an instance of VERTEX-COVER with a graph $G=(V,E)$ and an integer $k$, we construct an instance of HORN-MINTRUST with a set of assumptions $\Gamma$, a conclusion $c$, a cost function $C$, and an integer $k$. Let $\Gamma$ contain an assumption $v_{i}$ for each vertex $v_{i}\in V$ with cost 1, an assumption $v_{i}\implies e_{j}$ for each vertex $v_{i}\in V$ that is incident to edge $e_{j}\in E$ with cost 0, and an assumption $(e_{1}\land\dots\land e_{|E|})\implies c$ with cost 0. By construction, $\Delta\models c$ if and only if $\Delta\models e_{j}$ for all $e_{j}\in E$ and $((e_{1}\land\dots\land e_{n})\implies c)\in\Delta$. Furthermore, $\Delta\models e_{j}$ for all $e_{j}\in E$ if and only if $v_{i}\in\Delta$ and $(v_{i}\implies e_{j})\in\Delta$ for all $e_{j}\in E$. Since $(v_{i}\implies e_{j})\in\Gamma$ if and only if $v_{i}$ is incident to $e_{j}$, $\Delta\models e_{j}$ for all $e_{j}\in E$ if and only if $\Delta$ is a subset of the $v_{i}$s that corresponds to a vertex cover of $G$. The cost associated with $\Delta$ is equal to the size of the vertex cover, which is at most $k$.
HORN-MINTRUST is in NP. We construct a verifier for HORN-MINTRUST. On input a candidate solution $\Delta$, the verifier can check that the sum of the costs of the assumptions in $\Delta$ is less than or equal to $k$ in one pass. Since $\Delta\models c$ holds if and only if the conjunction of the Horn clauses in $\Delta$ with $(c\implies\bot)$ is unsatisfiable, and Horn-satisfiability is in P, checking that $\Delta\models c$ is possible in polynomial time. ∎

### v2

@@ 3.1. Problem definition
Given as input (1) a set of propositional formulas $\Gamma=\{A_{1},\ldots,A_{n}\}$ representing assumptions, (2) a propositional variable $c$ representing a conclusion, and (3) a function $C\colon\Gamma\to\mathbb{N}$ that assigns a cost to making each assumption, output a set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of assumptions in $\Delta$ is minimized, subject to the constraint $\Delta\models c$.
We provide a simple example of the minimum trust problem, continuing the example from Section 2. Suppose that aho_corasick depends on memchr, where aho_corasick was developed by Alice, and memchr was developed by Bob and has no dependencies. Table 1 gives a set of candidate assumptions $A_{1},\ldots,A_{8}$ a user can make about the safety of the aho_corasick library and their associated costs in this scenario. We are asked to find a subset of $A_{i}$ with minimum cost which entails $c$. In this example, the minimum cost is $45$, incurred from assumptions $a$ (cost 20), $b$ (cost $5$), $(a\land m)\implies c$ (cost 10), and $b\implies m$ (cost $10$). Note that the set of assumptions is not necessarily unique in general (as there may be a tie for the minimum cost), though it is unique in this example.
@@
Table 1. Example instance of the minimum trust problem.
Candidate assumptions, their encodings, and costs for the running example.
@@ 3.2. Naive algorithm
Algorithm 1 shows a naive solution for the minimum trust problem using an SMT solver by encoding the constraints as sentences over linear integer arithmetic and using quantifiers over propositional variables. To associate each assumption being made to a cost, we use auxiliary propositional variables $a_{1},\ldots,a_{n}$ to track each assumption $A_{1},\ldots,A_{n}$. Each $a_{i}$ is assigned to true if and only if $A_{i}$ is included in the solution set of assumptions $\Delta$. In the running example, we would have $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, etc. We then encode the cost of the assumptions in $\Delta$ as a sum of if-then-else (iteite) terms; in the running example, the cost of any particular solution is given by $\text{{{cost}}}(a_{1},\ldots,a_{8})=ite(a_{1},100,0)+\dots+ite(a_{8},10,0)$. The constraint that the conclusion $c$ follows from the assumptions being made, $\Delta\models c$, is encoded as a formula $\text{{{valid}}}(a_{1},\ldots,a_{n})$. In the running example, $\text{{{valid}}}(a_{1},\ldots,a_{8})=\neg\exists c,m,d,a,b\>((a_{1}\implies c)\wedge\dots\wedge(a_{8}\implies(b\implies m))\wedge\neg c)$. Finally, we encode using quantifiers the fact that $\text{{{cost}}}(a_{1},\ldots,a_{n})$ is minimized subject to the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$. In the running example, the SMT solver would produce a model that assigns $a_{5},a_{6},a_{7},a_{8}$ to true and $a_{1},a_{2},a_{3},a_{4}$ to false, incurring a cost of $45$.
@@
7:    $\Gamma\leftarrow\text{{{get\_all\_assumptions(lib)}}}$
@@
9:     $F\leftarrow F\wedge(a_{i_{\text{{{lib}}}}}\implies A_{i})$
@@
11:     $F\leftarrow F\wedge\text{{{JoinAssumptions}}}(\text{{{dep}}})$ $\triangleright$ JoinAssumptions is called recursively on each dependency
@@
13:  procedure SolveMinimumTrust()
14:    $F(a_{1},\ldots,a_{n})\leftarrow\text{{{JoinAssumptions(library)}}}$
15:    $\{v_{1},\ldots,v_{m}\}\leftarrow\text{{{atoms}}}(F)\setminus\{a_{1},\ldots,a_{n}\}$ $\triangleright$ atoms returns the set of propositional variables in a formula
16:    $\text{{{valid}}}(a_{1},\ldots,a_{n})\leftarrow\neg\exists c\exists v_{1}\ldots\exists v_{m}\>(F(a_{1},\ldots,a_{n})\wedge\neg c)$ $\triangleright$ $c$ is the conclusion ‘the library is safe’
17:    $\text{{{cost}}}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
18:    $\text{{{minimal}}}\leftarrow\forall a_{1}\ldots\forall a_{n}\>(\text{{{valid}}}(a_{1},\ldots,a_{n})\implies\text{{{cost}}}(a^{*}_{1},\ldots,a^{*}_{n})\leq\text{{{cost}}}(a_{1},\ldots,a_{n}))$
19:    $\text{{{S.solve}}}(\text{{{valid}}}(a^{*}_{1},\ldots,a^{*}_{n})\land\text{{{minimal}}})$ $\triangleright$ $(a^{*}_{1},\ldots,a^{*}_{n})$ is the solution to the minimum trust problem
@@
Algorithm 1 recursively calls JoinAssumptions on each library in the dependency tree on line 11 to construct the formula $F(a_{1},\ldots,a_{n})$ that represents the assumptions being made, then uses the above encoding to create the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$. The algorithm encodes that the cost of assumptions is minimized subject to the constraint $\text{{{valid}}}(a_{1},\ldots,a_{n})$ on line 18. ∎
Although it is correct, the use of quantifiers makes the first algorithm computationally expensive; it generates a linear number of nested quantifiers in the size of the dependency tree, which have to be expanded before applying a SAT solver. While some of this complexity is unavoidable (see Section 3.4), to address this we next propose an algorithm that does not use quantifiers, and which we have found to be more effective in practice.
@@ 3.3. Horn clause algorithm
The key idea of Algorithm 2 is to restrict the set of assumptions to be acyclic definite Horn clauses; in this case, the problem can be solved efficiently with a binary search reduction using quantifier-free queries to an SMT solver. The set of assumptions that can be expressed as acyclic definite Horn clauses is sufficient for all of the assumptions we identify based on prior work (Table 2(a)).
In the running example, observe that the set of conjuncts of $F$ that establish the safety of the aho_corasick library ($c$) are $a_{1}\implies c$, $a_{4}\implies(d\implies c)$, and $a_{7}\implies((a\land m)\implies c)$. Rewritten as Horn clauses with a body and a head, these conjuncts are $a_{1}\implies c$, $(a_{4}\land d)\implies c$, and $(a_{7}\land a\land m)\implies c$. Therefore, with the goal of showing that $c$ is true, the formula $a_{1}\lor(a_{4}\land d)\lor(a_{7}\land a\land m)$ must evaluate to true. The only conjunct of $F$ that establishes $d$ is $a_{3}\implies d$. Therefore, we can substitute $d$ with $a_{3}$ in the formula, resulting in $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a\land m)$. By repeating this unfolding process, we can simplify the formula to $a_{1}\lor(a_{4}\land a_{3})\lor(a_{7}\land a_{5}\land a_{6}\land a_{8})$. This formula will evaluate to true if and only if the assumptions made entail the conclusion $c$. Algorithm 2 uses this observation to solve the minimum trust problem when the assumptions are acyclic definite Horn clauses. Due to binary search, Algorithm 2 produces a logarithmic number of calls to the SMT solver with respect to the maximum cost, and each call is quantifier-free.
@@
5:    $F\leftarrow\text{{{simplify}}}(F)$ $\triangleright$ simplify the formula into a conjunction of Horn clauses
6:    $\Pi\leftarrow\text{{{clauses}}}(F)$ $\triangleright$ get the set of Horn clauses in $F$
@@
9:     if $\text{{{head}}}(\pi_{i})$ is the conclusion $c$ then
10:       $E\leftarrow E\lor\text{{{body}}}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $c$
11:    $\leftarrow\text{{{atoms}}}(F)\setminus\{a_{1},\ldots,a_{n}\} \triangleright VV$ stores the set of propositional variables in $F$ other than the $a_{i}$s
@@
15:       if $\text{{{head}}}(\pi_{i})$ is $v_{i}$ then
16:        $T\leftarrow T\lor\text{{{body}}}(\pi_{i})$ $\triangleright$ collect all the bodies of the Horn clauses with head $v_{i}$
@@
21:  procedure SolveHornMinimumTrust()
22:    $F(a_{1},\ldots,a_{n})\leftarrow\text{{{JoinAssumptions(library)}}}$ $\triangleright$ same as in Algorithm 1
23:    $\text{{{valid}}}(a_{1},\ldots,a_{n})\leftarrow\text{{{Unfold}}}(F)$
24:    $\text{{{cost}}}(a_{1},\ldots,a_{n})\leftarrow\sum_{i=1}^{n}ite(a_{i},C(a_{i}),0)$
25:    $\text{{{left}}}\leftarrow 0$ $\triangleright$ initialize the min cost as 0
26:    $\text{{{right}}}\leftarrow\sum_{i=1}^{n}C(a_{i})$ $\triangleright$ the max cost is the sum of the costs of all assumptions
27:    while $\text{{{left}}}<\text{{{right}}}$ do
28:     $k\leftarrow\lfloor(\text{{{left}}}+\text{{{right}}}+1)/2\rfloor$ $\triangleright$ binary search for min cost
29:     $\text{{{cost\_leq\_k}}}\leftarrow\text{{{S.solve}}}(\text{{{valid}}}(a_{1},\ldots,a_{n})\wedge\text{{{cost}}}(a_{1},\ldots,a_{n})\leq k)$
30:     if cost_leq_k is SAT then
31:       $\text{{{left}}}\leftarrow k$
@@
33:       $\text{{{right}}}\leftarrow k-1$
34:    $\text{{{min\_cost}}}\leftarrow\text{{{left}}}$
@@
Algorithm 2 is correct for assumptions consisting of acyclic definite Horn clauses.
@@
Algorithm 2 first constructs $E$ as the disjunction of all the bodies of the Horn clauses whose heads are the conclusion $c$ (line 10). The conclusion is provable from the assumptions if and only if $E$ evaluates to true. Algorithm 2 then proceeds to iterate through $V$ with the goal of eliminating all the propositional variables in $V$ from $E$. For each propositional variable $v_{i}$ in $V$, the algorithm constructs a formula $T$ consisting of the disjunction of the bodies of the Horn clauses whose heads are $v_{i}$ and then eliminates $v_{i}$ by substituting $T$ in place of $v_{i}$ in $E$, similar to the construction of $E$ on line 10. After the substitution, $E$ will still be a necessary and sufficient condition for the conclusion to be true. The substitution is then applied to each remaining clause in line 19; since the Horn clauses are acyclic, eliminating $v_{i}$ is guaranteed to remove $v_{i}$ from all future substitutions. This process is repeated until all propositional variables in $V$ are eliminated from $E$. Each step in Unfold maintains that $E$ is necessary and sufficient for the conclusion to be true. The formula generated by Unfold is used in a binary search to repeatedly query S for valid minimum trust solutions with different costs until the minimum cost is found. ∎
It should be noted that the unfolding procedure in Algorithm 2 is similar to the well-known SLD resolution algorithm for logic programming (Kowalski, 1974). Unfortunately, the generated formula has an exponential number of clauses in the worst case (with respect to the number of crates in the dependency tree). This formula, however, is in disjunctive normal form (DNF) after minor simplification, which means the SMT solver can solve each clause independently. Note that it is also possible to have a reduction that produces a polynomial-size formula, though this is not the reduction we use in our implementation.
@@ 3.4. Complexity
To establish the complexity of the minimum trust problem (Definition 1), we study the corresponding decision problem. To that end, define MINTRUST to be the problem that asks, given a tuple $(\Gamma,c,C,k)$ — consisting of a set of assumptions $\Gamma$, a conclusion $c$, a cost function $C$, and integer $k$ — does there exists a set of assumptions $\Delta\subseteq\Gamma$ such that the sum of the costs of the assumptions in $\Delta$ is less than or equal to $k$ and $\Delta\models c$? Similarly, define HORN-MINTRUST to be the restriction of this problem where $\Gamma$ consists of Horn clauses. The following theorems establish that the unrestricted minimum trust problem is computationally infeasible (Theorem 4) and that the Horn-restricted minimum trust problem is intractable but still within the reach of SMT solvers (Theorem 5).
@@
We reduce from Shortest Implicant Core (SIC), which is known to be $\Sigma_{2}^{P}$-hard (Schaefer and Umans, 2002) (Umans, 2001). The SIC problem asks: given a DNF formula $\varphi$, an implicant $I$ of $\varphi$ (a set of literals whose conjunction implies $\varphi$), and an integer $k$, does an implicant $I^{\prime}\subseteq I$ of size at most $k$ exist? Given an instance of SIC with a DNF formula $\varphi$, an implicant $I$ of $\varphi$, and an integer $k$, we construct a MINTRUST instance $(\Gamma,c,C,k)$. Let $\Gamma$ contain an assumption for each literal in $I$ with a cost of $1$ and the assumption $\varphi\implies c$ with a cost of 0.
By construction, $\Delta\models c$ if and only if $\Delta\models\varphi$ and $(\varphi\implies c)\in\Delta$. Furthermore, $\Delta\models\varphi$ if and only if $\Delta$ contains an implicant $I^{\prime}\subseteq I$ of $\varphi$. Then, the cost associated with $\Delta$ is equal to the size of $I^{\prime}$, so $\varphi$ has an implicant of size at most $k$ if and only if the MINTRUST instance has a solution of cost at most $k$. Therefore, an implicant $I^{\prime}\subseteq I$ of $\varphi$ of size at most $k$ exists if and only if this MINTRUST instance has a solution of cost at most $k$. ∎
@@
HORN-MINTRUST is NP-hard. We reduce from Vertex Cover (Karp, 1972). Given an instance of vertex cover with a graph $G=(V,E)$ and an integer $k$, we construct a HORN-MINTRUST instance $(\Gamma,c,C,k)$. Let $\Gamma$ contain an assumption $v_{i}$ for each vertex $v_{i}\in V$ with cost $1$, an assumption $v_{i}\implies e_{j}$ for each vertex $v_{i}\in V$ that is incident to edge $e_{j}\in E$ with cost 0, and an assumption $(e_{1}\land\dots\land e_{|E|})\implies c$ with cost 0. By construction, $\Delta\models c$ if and only if $\Delta\models e_{j}$ for all $e_{j}\in E$ and $((e_{1}\land\dots\land e_{|E|})\implies c)\in\Delta$. Furthermore, $\Delta\models e_{j}$ for each $e_{j}\in E$ if and only if $v_{i}\in\Delta$ and $(v_{i}\implies e_{j})\in\Delta$ for some incident $v_{i}$. It follows that $\Delta\models e_{j}$ for all $e_{j}\in E$ if and only if $\Delta$ is a subset of the $v_{i}$s for a vertex cover of $G$, together with the assumptions for incident edges. The cost associated with $\Delta$ is equal to the size of the vertex cover, which is at most $k$.
HORN-MINTRUST is in NP. We construct a verifier for HORN-MINTRUST. On input a candidate solution $\Delta$, the verifier checks that the sum of the costs of the assumptions in $\Delta$ is at most $k$ in one pass. Since $\Delta\models c$ holds if and only if the conjunction of the Horn clauses in $\Delta$ with $(c\implies\bot)$ is unsatisfiable, and Horn-satisfiability is in P, checking that $\Delta\models c$ is possible in polynomial time. ∎